# Bayesian model-data comparison incorporating theoretical uncertainties

### v2

@@ Appendix A Corner plots for model parameters
Figure 5: Corner plot of all model parameters (with median ±$\pm 68\%$ CI displayed on the diagonals) for the five-parameter model with parametrized $\eta/s$, using mock data generated with fixed $\eta/s=0.1$ and $\epsilon_{\rm sw}=0.2\,$GeV/fm3, corresponding to the results shown in Fig. 3 in the main text.
@@
Figure 6: Corner plot of all model parameters (with median ±$\pm 68\%$ CI displayed on the diagonals) for the five-parameter model with parametrized $\eta/s$, using mock data generated via the parametrization with $T_{\rm kink}=0.18\,{\rm GeV}$, $a_{\rm low}=-1$, $a_{\rm high}=1$, and $(\eta/s)_{\rm kink}=0.1$, along with $\epsilon_{\rm sw}=0.2\,$GeV/fm3. This plot corresponds to the results shown in Fig. 4 in the main text.

### v3

@@ Appendix A Corner plots of model parameters for heavy-ion simulation
Figure 6: Corner plot of all model parameters (with median $\pm 68\%$ CI displayed on the diagonals) for the five-parameter model with parametrized $\eta/s$, using mock data generated with fixed $\eta/s=0.1$ and $\epsilon_{\rm sw}=0.2\,$GeV/fm3, corresponding to the results shown in Fig. 4 in the main text.
@@
Figure 7: Corner plot of all model parameters (with median $\pm 68\%$ CI displayed on the diagonals) for the five-parameter model with parametrized $\eta/s$, using mock data generated via the parametrization with $T_{\rm kink}=0.18\,{\rm GeV}$, $a_{\rm low}=-1$, $a_{\rm high}=1$, and $(\eta/s)_{\rm kink}=0.1$, along with $\epsilon_{\rm sw}=0.2\,$GeV/fm3. This plot corresponds to the results shown in Fig. 5 in the main text.